Carlos trows a rock from the pier. The height of the rock h(t) can be moxled by the function. h(t)=-5t^2+13t+6. When will the rock hit the water ? What is the maximum height of the rock ?

Answer by stanbon(75887)

Carlos trows a rock from the pier. The height of the rock h(t) can be moxled by the function. h(t)=-5t^2+13t+6. When will the rock hit the water ? What is the maximum height of the rock ?
----
Its height is zero when the rock hits the water.
----
Solve -5t^2 + 13t + 6 = 0
Factor:
-5t^2 + 15t-2t + 6 = 0
-5t(t-3)-2(t-3) = 0
(t-3)(-5t-2) = 0
t = 3 seconds (time at which the rock hits the water
---
max height occurs when t = -b/(2a) = -13/(2*-5) = 13/10 = 1.3
---
Max height = h(1.3) = -5(1.3)^2 + 13(1.3) + 6 = -5*1.69 +16.9+6 = 14.49 ft
===================================
Cheers,
Stan H.
===================================